OpenStudy (anonymous):

Graph the linear by finding intercepts if possible. 4x+5y= 20

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=4x%2B5y%3D+20

Directrix (directrix):

4x+5y= 20 5y = 20 -4x y = 4 -(4/5)x y-intercept has coordinates (0,4) To get the x-intercept, y has to be 0. 5(0) = 20 - 4x 0 = 20 -4x -20 = -4x 5 = x x-intercept has coordinate (5,0) Plot (5,0) and ((0,4) and the two points (intercepts) will determine the line.